1. What do “3” and “p” imply in 3p? How many orbitals (ml values) are there in a 3p sub-level (draw them and name them)? What are the possible values ​​of the quantum spin number of the electron?

2. Give the values ​​of the quantum numbers associated with the following orbitals: a) 2p2, b) 3s2, c) 5d8.

3. Point out which of the following sets of quantum numbers are unacceptable and explain why:
    a) (1, 0, 1/2, +1/2), b) (3, 0, 0, +1/2), c) (2, 2, 1, + 1/2), d) ( 4, 3, -2, + 1/2), e) (3, 2, 1, 1).

4. Enter the electronic configuration of:
    a) The phosphorus atom present in phosphorous acid.
    b) The chlorine atom in iron chlorate III.
    c) The manganese atom in potassium permanganate.
Answer #1

To address any election in any atom, there are four Quantam number 4 principle Quantum no. ( n u r b It identifies the shellthese values are: - ma-p, -(4-1),-(4-2), --.0, 1 --(4-2) (4-1), f. for example, if we consider t=2 then possible values of msed and, f. Orbital name No. of orbital uleo 7. So 3 in the 3p sub-level, there orbitul , because p has 3 are orbitals. 22222m = -2 ott 2px 2py 2 P2 Quantum no. for 1st electrom, n= a f = 1 (for P, d = 1) ma-1 S=+1/2 (spin up) Quantum no. for and eleQuantum Mo for and election, n=3 f = 0 m=o s=-1/2 (spin down) c) 508 m = -2 -1 oth +2 possible quantum Mos. n = 5 (for all elSo, (1,0 ,f is not a possible Quantum . b) (3,0,0, +4/2) n=3 (possible = 0 (Corresponding to s) (possible) m=0 (possible) S =S=+4₂ (spin up) So, (4, 3, 2, +4/) is a possible quantum no e) ( 3, 2, 1, 1) n = 3 (possible) 1=2 (possible) mal possible) s=acid = 1s?, ?s?aps, as? b) iron (1) chlorate - PelclO3)₂ th Atomic no. of cd is L7 So, Ground state electronic configuration